OpenStudy (anonymous):

y=3x-1 4,1 2,5 4,3 0,-3

OpenStudy (zpupster):

here just plug in your x values and see which is true for the y value

HanAkoSolo (jamierox4ev3r):

right! let me the first one, just as an example. 1=3(4)-1 . That right translates to 1=12-1 Finally, that says that 1=11. However, we know that is not true so we try the rest until we find one that is correct. Make sense? @wolowizzard ?